OpenStudy (anonymous):

At the start of the basketball game, the referee tosses a ball for a jump ball. The equation that models the height of the ball is h(t) = -16t2 + 24t + 5 During what time interval is the height of the ball above 9 feet? (Round your answer to the nearest tenth of a second.)

OpenStudy (anonymous):

solve \[t(t)=-16t^2+24t+5=9\] for t. get to answers because it is a quadratic. it will be greater than 9 between them

OpenStudy (anonymous):

solutions are \[t=\frac{3\pm\sqrt{5}}{4}\] so it will be great than 9 on the interval \[\frac{3-\sqrt{5}}{4}<t<\frac{3+\sqrt{5}}{4}\]